OpenStudy (anonymous):

Find the LCM of 8, 12, and 16. A. 48 B. 96 C. 192 D. 1536

OpenStudy (anonymous):

only draw if your good at openstudies computer drawing

OpenStudy (cwrw238):

you can find LCM's by first finding the prime factors:- 8 = 2*2*2 12= 2*2*3 16 = 2*2*2*2

OpenStudy (anonymous):

so pretty much like this?

OpenStudy (anonymous):

^^^^^^^^^^^example

OpenStudy (cwrw238):

note that 2*2 is common to all three so just take 2 then multiply by all the other factors so LCM = 2*2*2*2*3 = 48

OpenStudy (cwrw238):

yes that's a good method to find the prime factors

OpenStudy (anonymous):

so pretty much like this?

OpenStudy (anonymous):

in crackly voice do u know who i am?

OpenStudy (cwrw238):

i'm not sure what you did in your last post....

OpenStudy (anonymous):

you said

OpenStudy (anonymous):

lcm = 2*2*2*2*3=48 so i did lcm is equal to 24*2 pretty much i guess which is 48

OpenStudy (cwrw238):

oh ok

OpenStudy (cwrw238):

yes 48 is the answer

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

question did my profile pic change on your screen?